OpenStudy (anonymous):

write y=3(x-4)(x-5) in standard form

OpenStudy (anonymous):

17 or 20.

OpenStudy (anonymous):

What? @ duckdynastylover

OpenStudy (anonymous):

@duckdynastylover

OpenStudy (jack1):

y=3(x-4)(x-5) y = (3x - 12)(x-5) use FOIL from here to multiply out the brackets y = (3x - 12)(x-5) y = 3x(x-5) - 12(x-5) y =

OpenStudy (jack1):

y = 3x(x-5) - 12(x-5) y = 3x^2 - 15x - 12x - 60 add the like terms to get final answer

OpenStudy (jack1):

@Maymayk